The weight (x grams) of a rose is assumed to be distrubuted as N (5, 12.4). A rose is selected randomly .find the probability that the weight of randomly selected flower is less than 4 gram

1
Expert's answer
2022-03-18T12:27:37-0400

"P(X<4)=P(Z<\frac{4-5}{\sqrt{12.4}})=P(Z<-0.28)=0.3897."